OpenStudy (anonymous):

solve the system 2x-2y+z=-1 x+3y+4z=11 3x+0y+2z=5

OpenStudy (wolf1728):

A) 2x-2y+z=-1 B) x+3y+4z=11 C) 3x+0y+2z=5 Multiply equation C by -5/2 C) -7.5x -5z = -12.5 Then add this to the other 2 equations A) 2x -2y +z= -1 B) x+ 3y +4z=11 which sums to D) -4.5x +y = -2.5 Multiply equation A by -6 A) -12x +12y -6z = 6 Then adding this to the other 2 equations B) x +3y +4z =11 C) 3x +0y +2z= 5 which sums to E) -8x +15y = 22 Dividing equation E by -15 (8/15)x -y = -22/15 Adding this to equation D D) -4.5x +y = -2.5 which sums to -3.9666666667 x = -3.9666666667 -x = -1 x = 1 Do you think you could solve for y and z?

OpenStudy (solomonzelman):

3x+0y+2z=5 , plug 1 for x

OpenStudy (solomonzelman):

so z=1

OpenStudy (solomonzelman):

y=-1

OpenStudy (wolf1728):

SolomonZelman Your calculation of y is incorrect 3x+0y+2z=5 Since x =1 then 3 + 2z = 5 2z = 2 z = 1 x+3y+4z=11 1 + 3y +4 = 11 3y = 6 y = 2